# Enhancing Robustness of CLIP to Common Corruptions through Bimodal Test-Time Adaptation

### v1

@@ 4 BAT-CLIP
Projection matching between the visual and text features. We propose learning visual and text features that are domain-specific and mutually aware by jointly updating the encoders, as in Figure 3 (middle row), opposed to a unimodal update in TPT. From Eq. 1, the visual feature $v_{k,i}^{t}$ of the $k^{th}$ image in batch $x_{i}^{t}$ needs to have a high similarity with the text feature $z_{c}$ of class c for good alignment. To quantify this similarity, a possible direction is to project the visual feature onto the text feature i.e., compute the scalar projection — $v_{k,i}^{t}\cdot\hat{z}_{c}$, where $\hat{z}_{c}$ is the normalized text feature of $\hat{z}_{c}$ i.e., $\hat{z}_{c}=\frac{z_{c}}{||z_{c}||_{2}}$. Geometrically speaking, maximizing this projection leads to more similarity between the corresponding features and hence, a better alignment.
A possible approach would be to compute the projection for each class-specific visual feature in a test batch. However, within the batch, due to the image corruption, the pseudo-labels could be wrong/noisy. Instead, we propose modeling the projection of the class prototype with its corresponding text feature. A prototype is useful, in such a scenario, because it encompasses the entire class distribution without relying on individual visual features [39]. In particular, we compute a class prototype as,
@@
i.e., we compute the mean feature of all the support visual features constituting a class c, based on the computed pseudo-labels (Eqs. 1 and 2), where $\hat{y}$ refers to the pseudo-labels computed via Eq. 1 and $\bar{v}_{c}$ is the class prototype of class c. Based on the class prototype, the projection matching loss is,
@@
Inter-class separability between class prototypes. The projection matching loss introduced in Eq. 3 encourages the $f_{vis}$ and $f_{txt}$ encoders to produce domain-specific, well-aligned, and input-aware features via jointly updating the encoders. However, when TTA occurs at a batch level with image corruption, visual features within the batch could overlap, leading CLIP to poorly differentiate between classes. This ultimately hinders effective class separation. Hence, with a desire to obtain separation between visual features from different classes, as illustrated in Figure 3 (last row), we propose maximizing the distance between the class prototypes. Since the visual and text features align across modalities via Eq. 3, class separation, in addition, is needed for good generalization, robustness, and adaptation. So, we increase the cosine distance between the class prototypes and enhance the discriminative nature as,
@@
where $\phi_{v}$ and $\phi_{t}$ refer to the parameters of $f_{vis}$ and $f_{txt}$, respectively.
Motivation of choice of parameters. Through thorough experimentation, [50] demonstrates that fine-tuning LayerNorm parameters in attention-based models leads to yielding strong results on multimodal tasks. This also brings a reduction in trainable parameters. Updating normalization layers has also been deeply studied in tasks involving domain shifts [21]. Motivated by these works, we update the LayerNorm parameters of both CLIP encoders. This constitutes updating $\sim 0.044$% of all parameters. For every new task/corruption, we reset the model parameters following TENT [46] i.e., we perform a single domain TTA. We use a generic hand-crafted prompt template “a photo of a ¡CLS¿.” at all times, thus avoiding the need for a prompt search. In summary, our bimodal test-time adaptation approach jointly updates the LayerNorm parameters of the CLIP encoders that are optimized synergistically through loss components aware of the input domain, leading to a more robust multimodal learning process.
@@ 5 Experiments and Results
5.1 Online Test-Time Adaptation
Baselines. We compare our approach with zero-shot CLIP and other proposed online TTA methods using CLIP - TPT [38] and VTE [9]  [^2] We do not benchmark against WATT [29] since they update the model several times, for each prompt template, over a number of prompt templates.. For TPT, 63 random augmentations are generated for a single test image in the vision space. Based on a threshold, the marginal entropy of confident predictions is minimized to optimize the prompts. In the text space, VTE considers an ensemble of hand-crafted prompts for a test image, to get the final predictions (Eq. 1). Alongside this, they also consider multiple random augmentations of the test image, with frozen CLIP encoders. We also contrast our work against state-of-the-art TTA methods i.e., TENT [46], RoTTA [47], RPL [34], and SAR [28]. In essence, we adopt these approaches to use CLIP. Following the guidelines set by [9], we update only the vision encoder $f_{vis}$ for such a setup. The model predictions can then be computed as in Eq. 1. We mention the training details of each TTA method in the Supplementary.
Implementation Details. We query the text encoder $f_{txt}$ with a general and fixed hand-crafted prompt template “a photo of a ¡CLS¿.” for all the datasets, as motivated earlier. We optimize both the vision encoder $f_{vis}$ and text encoder $f_{txt}$. For CIFAR-10C, we use an AdamW optimizer at a learning rate of $10^{-3}$. Similarly, for CIFAR-100C and ImageNet-C, Adam and AdamW optimizers are respectively used, at a fixed learning rate of 5$\times$$10^{-4}$, with the model being reset after each task. The batch sizes are set to 200, 200, and 64 for the datasets, following various TTA benchmarks, at a corruption severity level of 5 for each task. As visual backbones, we report results on ViT-B/16 and ViT-B/32 [10], where all of our experiments are run on a single NVIDIA RTX A5000 GPU.
Method
@@ 5.1 Online Test-Time Adaptation
ViT-B/16
37.92 41.7 54.42 71.75 40.89 67.93 73.62 73.89 77.35 70.22 84.45 62.36 53.81 47.65 59.43 61.16
@@
Gain/Loss(%)
+18.71 +17.83 +1.53 +7.63 +9.38 +10.96 +6.57 +6.29 +5.35 +8.72 +3.27 +19.79 +9.62 +2.09 +4.00 +9.70
ViT-B/32
35.47 39.94 43.23 69.95 41.43 64.50 70.13 70.85 72.33 66.66 81.37 64.57 59.69 48.28 56.62 59.00
@@
47.59 50.18 53.15 71.39 53.86 67.92 72.90 76.37 76.30 70.78 83.27 61.07 69.00 58.57 61.14 64.90
Ours
Gain/Loss(%)
+4.80 +5.81 -0.61 +5.40 +0.18 +6.98 +2.89 +1.30 +2.80 +4.53 +2.21 +10.99 -1.59 -1.51 +0.15 +3.36
CIFAR-100C
ViT-B/16
19.64 21.40 25.26 42.54 20.03 43.17 47.95 48.35 49.74 41.57 57.02 34.58 29.15 23.96 32.43 35.79
TPT
17.95 19.51 27.13 43.53 20.08 42.65 48.63 49.11 49.48 42.14 57.35 33.26 31.13 27.59 32.75 36.15
VTE
17.96 18.72 28.17 40.38 19.60 39.50 45.33 48.24 46.87 40.73 55.31 30.04 32.47 30.35 31.45 35.01
@@
Gain/Loss(%)
+5.27 +6.33 +5.49 +6.58 +6.19 +5.32 +6.22 +3.24 +1.88 +6.24 +5.92 +10.63 +2.27 +2.03 +4.56 +5.94
ViT-B/32
16.23 17.83 17.57 39.07 17.63 38.55 43.81 42.32 43.46 39.71 50.32 29.34 28.74 22.85 29.42 31.79
@@
16.84 18.33 18.94 39.63 22.88 39.13 43.80 44.56 44.88 39.21 49.37 28.37 34.13 26.87 30.12 33.14
Ours
Gain/Loss(%)
+4.51 +6.38 +3.38 +6.63 +0.19 +5.51 +6.11 +2.67 +2.00 +4.77 +7.62 +9.18 +0.43 +0.86 +3.07 +4.46
ImageNet-C
ViT-B/16
11.18 12.54 12.04 23.36 15.18 24.50 22.58 32.32 29.88 35.88 54.18 17.20 12.72 30.96 33.26 24.51
TPT
8.48 9.46 10.20 23.98 15.16 25.10 24.00 33.94 32.12 37.08 55.64 16.54 13.68 34.06 33.58 24.87
VTE
9.18 10.76 10.78 24.72 14.30 24.36 25.24 35.38 32.46 38.16 55.56 16.14 14.26 38.72 33.98 25.60
@@
Gain/Loss(%)
+8.14 +8.84 +7.56 +1.86 +6.76 +5.78 +3.78 +1.10 -0.46 +2.82 +1.08 +8.94 +9.48 -1.05 +4.36 +5.12
ViT-B/32
12.88 13.04 12.90 24.42 11.86 22.72 20.20 25.70 25.84 30.28 50.54 17.32 18.96 32.20 29.12 23.20
TPT
12.04 12.64 12.52 25.38 12.28 22.68 20.78 26.36 26.64 30.78 51.02 16.50 19.90 33.62 30.62 23.58
VTE
11.96 12.32 13.44 25.06 11.70 22.58 22.40 27.38 27.02 32.28 51.52 16.84 19.94 34.80 32.82 24.14
Ours
Gain/Loss(%)
+3.96 +5.16 +2.66 -0.34 +8.62 +6.18 +2.84 +2.04 +0.16 +3.74 -1.34 +0.34 +7.74 +1.40 +2.60 +3.25
Table 2: Mean accuracy (%) on CIFAR-10C, CIFAR-100C, and ImageNet-C - TTA mean accuracy of the 15 corruptions (tasks) at a severity level of 5, using ViT-B/16 and ViT-B/32. + and - denote the absolute gain/loss w.r.t the next best performance. We contrast our results against zero-shot ViT-B/16 and -B/32, TPT [38], and VTE [9].
@@
Figure 4: BAT-CLIP yields more discriminative visual features that exhibit stronger alignment with their corresponding text features. The t-SNE [43] plots show visual ($\circ$) and text ($\bigstar$) features for CIFAR-10C, comparing zero-shot ViT-B/16 with our approach.
Results on CIFAR-10C, CIFAR-100C, and ImageNet-C. We present the TTA results in Table 2. Our method significantly improves the mean accuracy across all tasks and datasets for both visual backbones. Unlike TPT [38] and VTE [9], which process a single test image at a time and therefore incur substantial time, our approach operates at the batch level. This allows us to leverage the semantic relationships within the batch. By updating the parameters of $f_{vis}$ and $f_{txt}$ through the projection matching loss and enhancing the cosine distance between the class prototypes, our method effectively strengthens the alignment between visual and text features and learns discriminative visual features. In Figure 4, we show t-SNE [43] plots of visual and text features, of a few tasks, and compare them against zero-shot ViT-B/16 for CIFAR-10C. It highlights how our approach enhances alignment with text features, fosters better class separation, and forms more distinct clusters, leading to significant improvements compared to zero-shot CLIP. The Supplementary shows detailed t-SNE plots for CIFAR-10C and CIFAR-100C. Concerning compute time per task, on a difficult dataset like ImageNet-C, our method takes about 45 s, compared to 40 mins for TPT and 4 mins for VTE. This makes our approach deployment-friendly for quick classification with large improvements.
Dataset Backbone Zero-Shot TENT RoTTA RPL SAR Ours
CIFAR-10C ViT-B/16 61.16 62.06 61.89 61.52 67.37 73.85
ViT-B/32 59.00 56.35 59.57 56.51 65.13 68.26
CIFAR-100C ViT-B/16 35.79 37.96 36.12 38.47 41.19 42.09
ViT-B/32 31.79 31.61 31.84 30.00 37.81 37.60
ImageNet-C ViT-B/16 24.51 25.15 24.78 25.08 29.73 30.72
ViT-B/32 23.20 24.05 23.33 23.62 28.07 27.39
Table 3: Mean accuracy (%) of CLIP adopted for TTA approaches at a corruption severity level of 5, using ViT-B/16 and ViT-B/32.
Results against prior TTA methods. In Table 3, we compare our method with existing TTA approaches, all adopted for CLIP. We report mean accuracy across all tasks. While most TTA methods outperform zero-shot CLIP on average, our approach consistently matches or exceeds the performance of these methods across all datasets and visual backbones. Detailed per-task accuracies are provided in the Supplementary. With a ViT-B/16 backbone, TENT [46] increases the accuracy over zero-shot CLIP by 0.9%, 2.18%, and 0.64%, for the respective benchmark datasets. SAR [28] shows significant performance improvements across both vision backbones over prior TTA methods. Although it performs similarly to our method on most benchmarks, our approach outperforms SAR on CIFAR-10C, achieving improvements of 6.48% and 3.13% on the ViT-B/16 and ViT-B/32 backbones, respectively. While all the prior approaches operate solely in the vision space, our approach leverages CLIP’s joint vision-language feature space, resulting in superior TTA performance.

### v2

@@ 4 Proposed Methodology
Projection matching between the visual and text features. We propose learning visual and text features that are domain-specific and mutually aware by jointly updating the encoders, as in Figure 4 (middle row), opposed to a unimodal. From Eq. 1, the visual feature $v_{k,i}^{t}$ of the $k^{th}$ image in batch $x_{i}^{t}$ needs to have a high similarity with the text feature $z_{c}$ of class c for good alignment. To quantify this similarity, an approach is to project the visual feature onto the text feature i.e., compute the scalar projection — $v_{k,i}^{t}\cdot\hat{z}_{c}$, where $\hat{z}_{c}$ is the normalized text feature of $\hat{z}_{c}$ i.e., $\hat{z}_{c}=\frac{z_{c}}{||z_{c}||_{2}}$. Geometrically speaking, maximizing this projection leads to more similarity between the corresponding features and hence, a better alignment. However, computing such a projection could have limitations. Within the test batch, due to the image corruption, the pseudo-labels could be wrong/noisy. Instead, we propose modeling the projection of the class prototype with its corresponding text feature. A prototype is useful, in such a scenario, because it encompasses the entire class distribution without relying on individual visual features [46]. In particular, we compute a class prototype as,
@@
i.e., for the current test batch, we compute the mean feature of all the support visual features constituting a class c, based on the computed pseudo-labels (Eqs. 1 and 2), where $\hat{y}$ refers to the pseudo-labels computed via Eq. 1 and $\bar{v}_{c}$ is the class prototype of class c. Based on the class prototype, the projection matching loss is,
@@
Inter-class separability between class prototypes. The projection matching loss introduced in Eq. 3 encourages the $f_{vis}$ and $f_{txt}$ encoders to produce domain-specific, well-aligned, and input-aware features via jointly updating the encoders. However, when TTA occurs at a batch level with image corruption, visual features within the batch could overlap, leading CLIP to poorly differentiate between classes. This ultimately hinders effective class separation. Hence, with a desire to obtain separation between visual features from different classes, as illustrated in Figure 4 (last row), we propose maximizing the distance between the class prototypes. Since the visual and text features align across modalities via Eq. 3, class separation, in addition, is needed for good generalization, robustness, and adaptation. Therefore, we additionally incorporate a loss to increase the cosine distance between the class prototypes,
@@
where $\phi_{v}$ and $\phi_{t}$ refer to the LayerNorm parameters of $f_{vis}$ and $f_{txt}$, respectively.
Motivation of choice of parameters. Through thorough experimentation, Zhao et al. [58] demonstrate that fine-tuning LayerNorm (LN) parameters in attention-based models leads to yielding strong results on multimodal tasks. This also brings a reduction in trainable parameters. Updating normalization layers has also been deeply studied in tasks involving domain shifts [26]. Motivated by these works, we update the LayerNorm parameters of both CLIP encoders. This constitutes updating $\sim 0.044$% of all parameters. For every new task/corruption, we reset the model parameters following TENT [53] i.e., we perform a single domain TTA. In summary, our bimodal online test-time adaptation approach jointly updates the LayerNorm parameters of the CLIP encoders that are optimized synergistically through loss components aware of the input domain, leading to a more robust multimodal learning process.
@@ 5 Experiments and Results
Baselines. We compare our approach with zero-shot CLIP (ZS), state-of-the-art online TTA methods adopted for CLIP following [10] i.e., TENT [53], RoTTA [54], RPL [41], and SAR [35]. We update only the vision encoder $f_{vis}$ for such a setup. We benchmark against other online TTA methods using CLIP - TPT [45] and VTE [10]. Based on the discussion in Section 2, we adapt the two variants of WATT [36], for our setup, to WATT-P* and WATT-S*, while maintaining method consistency. For a “fairer” comparison, on each test batch, the model is adapted once using each of the suggested 8 prompt templates. Note that this setup is not online yet. In addition, CLIP parameters are reset after every task. We report the results on CIFAR-10C and CIFAR-100C only, based on their suggested settings. On ImageNet-C, we also report results against online StatA [55], based on their chosen hyperparameters. The model predictions can then be computed as in Eq. 1. We mention the details of each TTA method in the Supplementary.
Implementation Details. We query the text encoder $f_{txt}$ with a general prompt template “a photo of a $<$CLS$>$.” for all the datasets, as motivated earlier. We optimize both the vision encoder $f_{vis}$ and text encoder $f_{txt}$. For CIFAR-10C, we use an AdamW optimizer at a learning rate of $10^{-3}$. Similarly, for CIFAR-100C and ImageNet-C, Adam and AdamW optimizers are respectively used, at a fixed learning rate of 5$\times$$10^{-4}$, with the model being reset after each task. For fairness to all the baselines, the batch sizes are set to 200, 200, and 64 for the datasets, following various TTA benchmarks, at a corruption severity level of 5 for each task. For WATT-P* and WATT-S*, we use the same templates and learning rates as mentioned in [36]. For StatA, the class correlation in a batch is controlled by a Dirichlet distribution ($\gamma$). We report results for $\gamma$ = 0.1 (low correlation) and -1 (sequential). All experiments are run on a single NVIDIA RTX A5000 GPU using ViT-B/16 as the visual backbone. Results on ViT-B/32 are reported in the Supplementary.
Method Venue
@@
ZS
ICLR’21 37.92 41.7 54.42 71.75 40.89 67.93 73.62 73.89 77.35 70.22 84.45 62.36 53.81 47.65 59.43 61.16
TENT
ICLR’21 15.49 18.28 38.12 81.59 21.73 76.32 82.35 84.62 82.19 80.60 91.83 80.55 63.52 58.57 54.71 62.03
RoTTA
CVPR’23 39.17 42.90 55.41 72.18 41.32 68.02 74.01 74.38 78.01 70.80 84.80 63.19 54.62 49.32 60.15 61.89
RPL
arXiv 15.47 17.43 40.73 81.76 20.08 69.89 82.93 84.43 83.19 81.84 91.80 79.42 64.89 54.07 54.90 61.52
SAR
ICML’22 47.98 53.60 60.56 74.30 47.56 73.15 76.43 77.91 79.88 75.66 86.79 71.62 58.34 62.03 64.71 67.37
[…]
@@
CIFAR-100C
ZS
ICLR’21 19.64 21.40 25.26 42.54 20.03 43.17 47.95 48.35 49.74 41.57 57.02 34.58 29.15 23.96 32.43 35.79
TENT
ICLR’21 7.60 8.21 8.33 51.81 7.95 52.45 55.34 54.16 36.17 50.92 65.63 54.51 36.52 43.99 35.81 37.96
RoTTA
CVPR’23 20.65 22.22 26.17 42.48 20.26 42.90 47.88 48.75 49.92 41.86 57.00 34.52 29.27 25.08 32.88 36.12
RPL
arXiv 6.44 7.09 7.09 52.16 11.81 52.33 55.50 54.20 38.83 51.99 66.07 54.45 36.86 42.83 39.45 38.47
SAR
ICML’22 25.30 27.19 32.78 47.12 23.42 47.16 51.70 51.94 52.48 48.77 61.54 44.50 32.26 33.67 38.06 41.19
[…]
@@
ECCV-W’24 17.96 18.72 28.17 40.38 19.60 39.50 45.33 48.24 46.87 40.73 55.31 30.04 32.47 30.35 31.45 35.01
WATT-P*
NeurIPS’24 20.53 22.22 27.3 43.14 17.51 42.37 48.17 47.31 49.34 41.49 57.07 35.29 27.75 25.83 31.89 35.81
WATT-S*
NeurIPS’24 21.20 23.11 28.23 44.16 18.45 43.44 49.14 48.16 50.05 42.35 57.82 36.43 28.36 26.85 32.93 36.71
@@
ImageNet-C
ZS
ICLR’21 11.18 12.54 12.04 23.36 15.18 24.50 22.58 32.32 29.88 35.88 54.18 17.20 12.72 30.96 33.26 24.51
TENT
ICLR’21 5.14 5.70 7.44 25.22 19.34 26.80 24.16 33.56 30.42 37.74 54.24 22.50 13.90 35.02 36.08 25.15
RoTTA
CVPR’23 11.34 12.96 12.32 23.38 15.50 24.66 22.90 32.56 30.02 35.98 54.32 17.20 12.80 31.06 33.46 24.78
RPL
arXiv 9.04 10.04 10.96 24.40 17.40 26.28 23.76 32.70 30.62 36.64 54.04 19.38 13.24 33.14 34.60 25.08
SAR
ICML’22 17.96 20.46 20.68 25.72 23.04 29.52 26.04 34.92 32.74 39.00 55.00 27.14 19.64 36.66 37.50 29.73
[…]
@@
ECCV-W’24 9.18 10.76 10.78 24.72 14.30 24.36 25.24 35.38 32.46 38.16 55.56 16.14 14.26 38.72 33.98 25.60
StatA ($\gamma$=0.1)
CVPR’25 10.56 11.22 10.86 22.11 14.12 22.39 20.85 31.39 29.90 34.63 53.32 16.00 12.46 29.55 32.79 23.47
StatA ($\gamma$=-1)
CVPR’25 11.83 13.02 12.41 23.71 15.02 23.64 21.79 31.80 30.22 36.52 54.15 17.56 13.00 32.01 33.37 24.67
@@
Table 1: Mean accuracy (%) on CIFAR-10C, CIFAR-100C, and ImageNet-C - TTA mean accuracy of the 15 corruptions (tasks) at a severity level of 5, using ViT-B/16.
5.1 Results: Online Test-Time Adaptation
Results on CIFAR-10C, CIFAR-100C, and ImageNet-C. We present the online TTA results in Table 1. While most TTA methods outperform zero-shot CLIP on average, BATCLIP consistently exceeds the performance of these methods across all datasets. SAR shows significant performance improvements over others. Although it performs similarly to our method on CIFAR-100C and ImageNet-C, BATCLIP outperforms SAR on CIFAR-10C, achieving improvements of 6.48%. While all the prior approaches operate solely in the vision space, our approach leverages CLIP’s joint vision-language feature space, resulting in superior TTA performance. For TTA approaches using CLIP, BATCLIP surpasses TPT and VTE across all tasks and datasets. For our WATT-P* and WATT-S* implementation, adaptation happens once on each prompt template for a batch. Interestingly, despite CLIP being trained on the batch by averaging adapted weights from multiple templates before inference, the performance remains subpar. This indicates that WATT [36] is not well suited for online TTA. In StatA [55], $\gamma$ controls the number of effective classes in a batch, for zero-shot VLM adaptation. On ImageNet-C, the large drop in performance for each task indicates the need for CLIP adaptation to severe image degradation and the importance of relevant text features. In Figure 5, we show t-SNE [49] plots of visual and text features, of a few tasks, and compare them against zero-shot ViT-B/16 for CIFAR-10C. It highlights how our approach enhances alignment with text features, fosters better class separation, and forms more distinct clusters, leading to significant improvements. The Supplementary shows detailed t-SNE plots for CIFAR-10C and CIFAR-100C.
@@ 5.1 Results: Online Test-Time Adaptation
Figure 5: BATCLIP yields more discriminative visual features that exhibit stronger alignment with their corresponding text features. The t-SNE [49] plots show visual ($\circ$) and text ($\bigstar$) features for CIFAR-10C, comparing zero-shot ViT-B/16 with our approach.
Dataset Domain ZS TENT WATT-P* WATT-S* Ours
OfficeHome Art 78.29 79.4 80.35 80.43 79.86
Clipart 64.03 63.18 66.85 66.9 66.44
Product 87.11 88.42 87.5 87.54 88.51
Real World 88.96 89.6 90.04 89.99 89.67
PACS Art 97.22 98.05 97.75 97.75 97.56
Cartoon 97.4 97.65 97.53 97.53 97.48
Photo 99.58 99.58 99.59 99.52 99.72
Sketch 86.23 88.75 88.52 88.65 87.76
VLCS Caltech101 99.43 99.43 99.36 99.36 99.51
LabelMe 68.15 68.14 66.92 68.49 68.94
SUN09 73.4 73.4 74.53 74.68 75.23
VOC2007 84.75 84.75 84.0 84.03 85.6
TerraInc L38 20.3 20.3 27.79 29.08 37.33
L43 31.52 31.52 33.98 34.13 32.71
L46 28.98 28.98 27.07 28.13 31.05
L100 52.35 52.35 43.59 42.32 55.22
Table 2: Accuracy on domain generalization datasets [14] with other domain shifts using a ViT-B/16 visual backbone.
5.2 Results: Domain Generalization Datasets
In this section, we evaluate our BATCLIP on commonly used domain generalization datasets [14] and benchmark against zero-shot ViT-B/16 (ZS), TENT, WATT-P* and WATT-S*. From Table 2, we observe that BATCLIP does better or comparable, across all datasets and domain shifts, to all the baselines. TENT matches zero-shot CLIP since LN normalization works per image across the features, independent of the batch statistics. We demonstrate that BATCLIP can be generalized and be robust to other various domain shifts, with a single step adaptation and global prompt template as opposed to WATT-P* and WATT-S* that use multiple prompt templates. The largest performance improvements come on the Terra Incognita dataset. The implementation details are described in the Supplementary.